OpenStudy (anonymous):

(4sqrt(7)-8sqrt(2)) (3sqrt(7)+6sqrt(2)) multiply

OpenStudy (anonymous):

-32sqrt14 * 18sqrt14 = -32 * 18 * 14 = -8064

OpenStudy (anonymous):

oh sorry it is -12 i took the - and + signs as multiplication by mistake